# Quaternion-Domain Super MDS for Robust 3D Localization

### v1

@@ IV-A2 Construction of the Quaternion-Domain GEK Matrix
To construct the quaternion-domain GEK matrix in (IV-A1), additional phase difference information is required—beyond the typically measurable pairwise distance $d_{m}$ and ADoA $\alpha_{mp}$—when the positional relationship between nodes is projected onto each plane. Depending on the extent to which this additional information can be obtained from measurements, two practical scenarios are considered.
@@
Figure 2: Parameters that can be obtained using a planar antenna.
Finally, we conclude this subsection by summarizing the QD-SMDS algorithm in the form of pseudo-code, presented in Algorithm 2 at the top of the next page.
Algorithm 2 Quaternion-Domain SMDS
1:Input:
3:The coordinates of at least 4 ANs.
4:Steps:
5:Construct the quaternion-domain GEK matrix $\tilde{\bm{K}}$ in (IV-A1) using the input parameters.
7:Obtain the edge vector $\hat{\bm{\nu}}$ using Eq. (33).
9:Compute $\hat{\bm{X}}$ from $\hat{\bm{V}}$ using Eq. (24).
@@ IV-C1 Simulation Conditions
Computer simulations were conducted to evaluate the performance of the proposed QD-SMDS algorithm. The simulation environment assumes a room with a dimensions of $30$[m] (length) $\times$ $30$[m] (width) $\times$ $10$[m] (height). ANs were placed at five locations: the four upper corners of the room, specifically at $(x,y,z)=(0,0,10),(30,0,10),(30,30,10),$ and $(0,30,10)$, as well as the origin $(x,y,z)=(0,0,0)$. TNs were randomly placed at 15 locations within the interior, with their $\mathrm{x},\mathrm{y}$, and $\mathrm{z}$ coordinates independently drawn from a uniform distribution.
@@
In turn, angle measurement errors $\delta_{\theta}$ are assumed to follow a Tikhonov-distribution [33, 34]. The PDF of the measured angle $\tilde{\theta}=\theta+\delta_{\theta}$ corresponding to the true angle $\theta$ is given by
$p_{\mathrm{\Theta}}\left(\tilde{\theta};\theta,\rho\right)=\frac{1}{2\pi I_{0}(\rho)}\exp\left[\rho\cos(\theta-\tilde{\theta})\right].$ (38)
@@
[…]
Estimation errors are evaluated using the Frobenius norm of the difference between the estimated TN coordinate matrix $\hat{\bm{X}}_{\mathrm{T}}$ and the true coordinate matrix $\bm{X}_{\mathrm{T}}$, i.e.,
$\xi\triangleq\frac{1}{N_{t}}\|\hat{\bm{X}}_{\mathrm{T}}-\bm{X}_{\mathrm{T}}\|_{\mathrm{F}}.$ (40)
Figure 3: Comparison of localization accuracy between the SMDS and QD-SMDS algorithms in Scenario I.
Figure 4: Comparison of localization accuracy between the SMDS and QD-SMDS algorithms in Scenario II.
@@ IV-C2 Simulation Results Without Missing Data
Based on the above simulation conditions, we have compared the localization accuracy of the conventional SMDS and the proposed QD-SMDS algorithms under both Scenario I and Scenario II described in Section IV-A. Fig. 3 shows the localization accuracy comparison between SMDS and QD-SMDS in Scenario I. The horizontal axis represents the standard deviation of the Gamma-distributed distance measurements, as defined in (37), while the vertical axis indicates the averaged estimation error, as defined in (40). The results are plotted for various angular measurement errors $\epsilon\in\left\{10^{\circ},20^{\circ},30^{\circ},40^{\circ},50^{\circ}\right\}$.
@@
When the angle error is small ($\epsilon=10^{\circ}$ and $20^{\circ}$), we can observe that the relative performance of the two methods varies depending on the distance error. The QD-SMDS algorithm outperforms the SMDS algorithm up to $\sigma_{d}=1.0$ [m] for $\epsilon=10^{\circ}$, and up to $\sigma_{d}=1.8$ [m] for $\epsilon=20^{\circ}$. However, as the distance error increases further, SMDS begins to achieve higher accuracy. This behavior can be attributed to the fact that when angle errors are minimal, the GEK matrix can be constructed with high precision, making aggressive noise suppression via low-rank truncation less critical. Furthermore, SMDS, which requires fewer (and noisy) parameters to construct the GEK matrix, becomes advantageous under such conditions.
@@
In contrast, when the angle error exceeds $30^{\circ}$, the QD-SMDS algorithm consistently outperforms SMDS, with the performance gap widening as the angle error increases. These results indicate that as the accuracy of the GEK matrix deteriorates, the role of SVD-based noise suppression through low-rank truncation becomes increasingly important. As such, QD-SMDS —which constructs a rank-$1$ GEK matrix—offers improved robustness to angle errors compared to SMDS, which employs a rank-$3$ GEK matrix.
@@
Fig. 4 shows the localization accuracy comparison between SMDS and QD-SMDS in Scenario II. As observed in Fig. 3, SMDS achieves higher localization accuracy when the angle error is up to $\epsilon=20^{\circ}$. However, beyond this point, QD-SMDS significantly outperforms SMDS, and the performance gap becomes more pronounced than in Scenario I. This improvement is attributed to the enhanced robustness of QD-SMDS against measurement errors, enabled by the additional angular information that can be measured or estimated using planar antennas.
IV-C3 Simulation Results With Missing Data
@@
We now turn our attention to a more practical scenario in which missing data results in a sparse GEK matrix. In Scenario I, since the SMDS algorithm is initially employed for data supplementation, an effective performance comparison involving matrix completion is not feasible; therefore, this subsection focuses on Scenario II. In our setup, $30\%$ of the entries in both the real-domain GEK matrix $\bm{K}_{\mathrm{r}}$ and the quaternion-domain GEK matrix $\bm{K}_{\mathrm{q}}$ are randomly removed, while preserving matrix symmetry. For SMDS, the incomplete real-domain GEK matrix is first completed using the low-rank matrix completion method described in [31], followed by execution of the SMDS algorithm. For QD-SMDS, the incomplete quaternion-domain GEK matrix is first decomposed into two complex matrices, each of which is individually completed using the same low-rank matrix completion method. These are then recombined into a single quaternion matrix, after which the QD-SMDS algorithm is executed.
Fig. 5 shows the localization accuracy comparison between SMDS and QD-SMDS in Scenario II under missing data conditions. As observed, when the angle error is small (e.g., $\epsilon=10^{\circ}$ and $20^{\circ}$), SMDS still yields the best performance, although the gap between SMDS and QD-SMDS is negligible. In contrast, as the angle error increases, QD-SMDS significantly outperforms SMDS, with the performance gap becoming even more pronounced than in the case without missing data. This improvement is attributed to the lower rank of the quaternion-domain GEK matrix compared to the real-domain one. Since low-rank structures are more favorable for matrix completion, QD-SMDS benefits from improved recovery accuracy, which directly enhances localization performance under larger angle errors.
@@ V Quaternion-Domain MRC-SMDS
In QD-SMDS, the primary source of computational complexity is the QSVD process. As described in Section II-B, QSVD requires performing SVD on a complex-valued matrix that is twice the size of the original quaternion-domain GEK matrix. To reduce this computational overhead, this section revisits the structural properties of the quaternion-domain GEK matrix and introduces low-complexity variants of the QD-SMDS algorithm that eliminate the need for QSVD.
@@ V-A QD-MRC-SMDS
where $\bm{\chi}_{\mathrm{A}}$ and $\bm{\chi}_{\mathrm{T}}$ denotes the quaternion coordinate vectors corresponding to the real-valued coordinate matrices $\bm{X}_{\mathrm{A}}$ and $\bm{X}_{\mathrm{T}}$ for ANs and TNs, respectively, as defined in (14) and (15).
The corresponding structure matrices are given by
@@
From (42), it is evident that the estimation of TN coordinates can be accomplished by accurately estimating the quaternion edge vector $\bm{\nu}_{\mathrm{AT}}$. Furthermore, based on (41), the quaternion-domain GEK matrix in (IV-A1) can be expressed as
@@
By exploiting the structure observed in (44), the edge vector can be estimated through simple linear filtering without the need for more complex operations.
First, using (42), $\bm{K}_{2}$ can be rewritten as
Finally, taking into account the non-commutativity of both matrix and quaternion multiplications, we solve (46) for $\bm{\chi}_{\mathrm{T}}$, thereby obtaining a closed-form expression for the estimated TN coordinates:
This low-complexity variant of the QD-SMDS algorithm is hereafter referred to as QD-MRC-SMDS, with its pseudo-code provided in Algorithm 3. Compared to the original QD-SMDS, QD-MRC-SMDS significantly reduces computational complexity by eliminating expensive operations such as QSVD, matrix inversion, and Procrustes transformation. Instead, it enables direct coordinate estimation via simple quaternion-domain multiplications by utilizing a specific subset of the quaternion-domain GEK matrix.
@@
[…]
@@ V-B Iterative QD-MRC-SMDS
Algorithm 4 Iterative QD-MRC-SMDS
1:Input:
3:The coordinates of all ANs.
4:Steps:
5:Construct $\bm{\nu}_{\mathrm{AA}}$, $\bm{K}_{2}$ and $\bm{K}_{3}$ via equations in Eqs. (28) and (IV-A1).
6:Initialize $\bm{\nu}_{\mathrm{AT}}$ using Eq. (50).
7:for $\tau=1,\ldots,\tau_{\mathrm{max}}$
8:Update $\bm{\nu}_{\mathrm{AT}}$ using Eq. (49).
9:end for
10:Obtain $\hat{\bm{\chi}}_{T}$ using Eq.(51).
This iterative variant of the QD-MRC-SMDS algorithm is hereafter referred to as iterative QD-MRC-SMDS, with its pseudo-code provided in Algorithm 4.

### v2

@@ IV-A2 Construction of the Quaternion-Domain GEK Matrix
Figure 2: Parameters that can be obtained using a planar antenna.
Algorithm 2 Quaternion-Domain SMDS
1:Input:
3:The coordinates of at least 4 ANs.
4:Steps:
5:Construct the quaternion-domain GEK matrix $\tilde{\bm{K}}$ in (32) using the input parameters.
7:Obtain the edge vector $\hat{\bm{\nu}}$ using Eq. (33).
9:Compute $\hat{\bm{X}}$ from $\hat{\bm{V}}$ using Eq. (24).
To construct the quaternion-domain GEK matrix in (32), additional phase difference information is required—beyond the typically measurable pairwise distance $d_{m}$ and ADoA $\alpha_{mp}$—when the positional relationship between nodes is projected onto each plane. Depending on the extent to which this additional information can be obtained from measurements, two practical scenarios are considered.
@@
Finally, this subsection concludes with a summary of the QD-SMDS algorithm in the form of pseudo-code, presented in Algorithm 2, together with a high-level architectural overview of the processing pipeline shown in Fig. 3.
[…]
Figure 3: High-level architecture of the proposed quaternion-domain localization pipeline.
@@ IV-C1 Simulation Conditions
Figure 4: Comparison of average estimation error between the SMDS and QD-SMDS algorithms in Scenario I.
Figure 5: Comparison of the empirical CDF between the SMDS and QD-SMDS algorithms in Scenario I.
Computer simulations were conducted to evaluate the performance of the proposed QD-SMDS algorithm. The simulation environment assumes a room with a dimensions of 3030m (length) $\times$ 3030m (width) $\times$ 1010m (height). ANs were placed at five locations: the four upper corners of the room, specifically at $(x,y,z)=(0,0,10),(30,0,10),(30,30,10),$ and $(0,30,10)$, as well as the origin $(x,y,z)=(0,0,0)$. TNs were randomly placed at 15 locations within the interior, with their $\mathrm{x},\mathrm{y}$, and $\mathrm{z}$ coordinates independently drawn from a uniform distribution.
@@
In turn, angle measurement errors $\delta_{\theta}$ are assumed to follow a Tikhonov-distribution [38, 39], with the PDF of the measured angle $\tilde{\theta}=\theta+\delta_{\theta}$ corresponding to the true angle $\theta$ is given by
[…]
@@
$\epsilon\triangleq\theta_{\mathrm{B}}\Big|\int_{-\theta_{\mathrm{B}}}^{\theta_{\mathrm{B}}}p_{\Theta}(\phi;0,\rho)d\phi=0.9.$ (39)
The following performance metrics are considered the evaluations hereafter:
  1. 1.
    Average estimation error, defined as
    $\xi\triangleq\frac{1}{L}\sum_{l=1}^{L}\frac{1}{N_{\mathrm{T}}}\left\|\hat{\bm{X}}_{\mathrm{T}}^{(l)}-\bm{X}_{\mathrm{T}}^{(l)}\right\|_{\mathrm{F}}=\frac{1}{L}\sum_{l=1}^{L}\xi^{(l)},$ (40)
    where $\hat{\bm{X}}_{\mathrm{T}}^{(l)}$ and $\bm{X}_{\mathrm{T}}^{(l)}$ denote the estimated and true coordinate matrices of the TNs, respectively, in the $l$-th Monte Carlo trial; $L$ is the total number of Monte Carlo trials; and $\xi^{(l)}\triangleq\frac{1}{N_{\mathrm{T}}}\big\|\hat{\bm{X}}_{\mathrm{T}}^{(l)}-\bm{X}_{\mathrm{T}}^{(l)}\big\|_{\mathrm{F}}$ represents the normalized Frobenius norm of the estimation error in the $l$-th trial.
  2. 2.
    Empirical CDF of the localization error, defined as
    $F_{E}(r)=\frac{1}{L}\sum_{l=1}^{L}\mathbf{1}\{\xi^{(l)}\leq r\},\quad r\geq 0,$ (41)
    where $\mathbf{1}\{\cdot\}$ is the indicator function.
@@ IV-C2 Simulation Results Without Missing Data
Figure 6: Comparison of average estimation error between the SMDS and QD-SMDS algorithms in Scenario II.
Figure 7: Comparison of empirical CDF between the SMDS and QD-SMDS algorithms in Scenario II.
@@
Based on the above simulation conditions, we have compared the localization accuracy of the conventional SMDS and the proposed QD-SMDS algorithms under both Scenario I and Scenario II described in Section IV-A. Fig. 4 and 5 show the localization accuracy comparison between SMDS and QD-SMDS in Scenario I. In Fig. 4, the horizontal axis represents the standard deviation of the Gamma-distributed distance measurements, as defined in (37), while the vertical axis indicates the averaged estimation error, as defined in (40). The results are plotted for various angular measurement errors $\epsilon\in\left\{10^{\circ},20^{\circ},30^{\circ},40^{\circ},50^{\circ}\right\}$. In Fig. 5, the horizontal axis represents the average estimation error, whereas the vertical axis shows the empirical CDF, i.e., the fraction of errors that do not exceed $\xi$ meters. The results are shown for two combinations of angular and distance measurement errors, $(\sigma_{d},\epsilon)=\{(1,20^{\circ}),(4,50^{\circ})\}$.
@@
When the angle error is small ($\epsilon=10^{\circ}$ and $20^{\circ}$), we can observe in Fig. 4 that the relative performance of the two methods in terms of the average estimation error varies depending on the distance error. The QD-SMDS algorithm outperforms the SMDS algorithm up to $\sigma_{d}=1.0$ m for $\epsilon=10^{\circ}$, and up to $\sigma_{d}=1.8$ m for $\epsilon=20^{\circ}$. However, as the distance error increases further, SMDS begins to achieve higher accuracy. This behavior can be attributed to the fact that when angle errors are minimal, the GEK matrix can be constructed with high precision, making aggressive noise suppression via low-rank truncation less critical. Furthermore, SMDS, which requires fewer (and noisy) parameters to construct the GEK matrix, becomes advantageous under such conditions. This limitation of QD-SMDS defines its operating regime, in which conventional SMDS alone may be sufficient to achieve high accuracy when the angular measurement errors are small.
@@
In contrast, when the angle error exceeds $30^{\circ}$, the QD-SMDS algorithm consistently outperforms SMDS, with the performance gap widening as the angle error increases. These results indicate that as the accuracy of the GEK matrix deteriorates, the role of SVD-based noise suppression through low-rank truncation becomes increasingly important.
As such, QD-SMDS —which constructs a rank-$1$ GEK matrix—offers improved robustness to angle errors compared to SMDS, which employs a rank-$3$ GEK matrix. On the other hand, Fig. 5 shows a clear advantage of QD-SMDS in terms of the empirical CDF: its distribution rises earlier than that of SMDS, indicating consistently higher accuracy across the entire error range for both combinations of angle and distance errors.
We conclude the discussion of Scenario I by noting that, although the practicality of this setup may appear limited due to the two-step procedure, similar approaches—where rough position estimates are first obtained to infer angular information and subsequently refined in a second step—have been considered in prior work [19, 21, 22]. Floating-point operation (FLOP) counts for each proposed algorithm are reported later in Table II. These results can assist users in determining whether the performance improvement achieved by this two-step procedure justifies the additional computational cost, given their available computational resources.
Fig. 6 compares the localization accuracy of SMDS and QD-SMDS in Scenario II in terms of the average estimation error. As also observed in Fig. 4, SMDS achieves higher localization accuracy when the angle error is up to $\epsilon=20^{\circ}$.
@@
However, beyond this threshold, QD-SMDS significantly outperforms SMDS, and the performance gap becomes more pronounced than in Scenario I. This improvement is attributed to the enhanced robustness of QD-SMDS against measurement errors, enabled by the additional angular information that can be measured or estimated using planar antennas.
A similar trend can be observed in Fig. 7. For small angular and distance errors, the SMDS curve rises earlier. In the second subfigure, corresponding to larger angular and distance errors, however, the pronounced leftward shift of the QD-SMDS CDF confirms a substantial reduction in localization error, demonstrating that the proposed method outperforms SMDS by markedly reducing large-error estimates.
IV-C3 Simulation Results With Random Missing Entries in GEK
We now turn our attention to a practical scenario in which missing data results in a sparse GEK matrix. In Scenario I, since the SMDS algorithm is initially employed for data supplementation, an effective performance comparison involving matrix completion is not feasible; therefore, this subsection focuses on Scenario II.
@@ IV-C3 Simulation Results With Random Missing Entries in GEK
Figure 9: Comparison of localization accuracy between SMDS and QD-SMDS with 10% missing distances and full angle information. The EDM matrix is completed using the algorithm proposed in [35].
In our setup, $30\%$ of the entries in both the real-domain GEK matrix $\bm{K}_{\mathrm{r}}$ and the quaternion-domain GEK matrix $\bm{K}_{\mathrm{q}}$ are randomly removed, while preserving matrix symmetry. For SMDS, the incomplete real-domain GEK matrix is first completed using the interest zone matrix approximation (IZMA) algorithm described in [36], followed by execution of the SMDS algorithm. The completion algorithm is terminated when both the reconstruction error on observed entries, and the update of the nuclear-norm threshold parameter $\lambda$ become smaller than 0.01. For QD-SMDS, the incomplete quaternion-domain GEK matrix is first decomposed into two complex matrices, each of which is individually completed using the same low-rank matrix completion method. These are then recombined into a single quaternion matrix, after which the QD-SMDS algorithm is executed.
Fig. 8 shows the localization accuracy comparison in terms of the average estimation error between SMDS and QD-SMDS in Scenario II under missing data conditions. As observed, when the angle error is small (e.g., $\epsilon=10^{\circ}$ and $20^{\circ}$), SMDS still yields the best performance, although the gap between SMDS and QD-SMDS is negligible. In contrast, as the angle error increases, QD-SMDS significantly outperforms SMDS, with the performance gap becoming even more pronounced than in the case without missing data.
This improvement is attributed to the lower rank of the quaternion-domain GEK matrix compared to the real-domain one. Since low-rank structures are more favorable for matrix completion, QD-SMDS benefits from improved recovery accuracy, which directly enhances localization performance under larger angle errors.
@@
IV-C4 Simulation Results with Missing Distances
A more challenging (and relevant) impairment occurs when distance measurements are unavailable due to the loss of line-of-sight between an anchor and a target. In this case, the degradation is more severe because a single missing distance induces a structured pattern of missing data in the GEK matrix: the corresponding row and column become entirely unobserved, resulting in 2M12M-1 missing GEK entries.
Under such conditions, directly applying IZMA to complete the GEK matrix is not viable. The completion method estimates missing entries by fitting coefficients (linear weights) from the observed elements. If an entire row (or column) is unobserved, the associated coefficients remain unconstrained and the recovery becomes ill-posed. Consequently, low-rank completion methods cannot, in general, uniquely recover fully unobserved rows/columns.
Instead, we first complete the EDM (i.e., the missing distances) and subsequently construct a complete GEK matrix to run SMDS and QD-SMDS. Specifically, we consider a setting similar to Scenario II, in which full angle information is available but 10% of the distances are missing. The missing distances are recovered using the semidefinite relaxation (SDR)-based EDM completion algorithm in [35] with $\lambda=1000$. The results in Fig. 9 exhibit the same qualitative trend as the case with randomly missing GEK entries. In particular, QD-SMDS begins to outperform SMDS only when the angular errors exceed $\epsilon=20^{\circ}$, for the same reasons discussed above.
@@ V Quaternion-Domain MRC-SMDS
In QD-SMDS, the primary source of computational complexity is the QSVD process. As described in Section II-B, QSVD requires performing an SVD on a complex-valued matrix whose size is twice that of the original quaternion-domain GEK matrix. To reduce this computational overhead, this section revisits the structural properties of the quaternion-domain GEK matrix and introduces low-complexity variants of the QD-SMDS algorithm that eliminate the need for QSVD.
@@ V-A QD-MRC-SMDS
where $\bm{\chi}_{\mathrm{A}}$ and $\bm{\chi}_{\mathrm{T}}$ denotes the quaternion coordinate vectors corresponding to the real-valued coordinate matrices $\bm{X}_{\mathrm{A}}$ and $\bm{X}_{\mathrm{T}}$ for ANs and TNs, respectively, as defined in (14) and (15). The corresponding structure matrices are given by
@@
From (43), it is evident that the estimation of the TN coordinates can be achieved by accurately estimating the quaternion edge vector $\bm{\nu}_{\mathrm{AT}}$. Furthermore, based on (42), the quaternion-domain GEK matrix in (32) can be expressed as
@@
7:Compute the estimated quaternion coordinate vector $\hat{\bm{\chi}}_{\mathrm{T}}$ using Eq. (48)
@@
By exploiting the structure observed in (45), the edge vector can be estimated using simple linear filtering, without the need for more complex operations.
First, using (43), $\bm{K}_{2}$ can be rewritten as
Finally, taking into account the non-commutativity of both matrix and quaternion multiplications, (47) is solved for $\bm{\chi}_{\mathrm{T}}$, thereby yielding a closed-form expression for the estimated TN coordinates [^3] Since the condition $\bm{B}_{\mathrm{AT}}^{\mathsf{T}}\bm{B}_{\mathrm{AT}}=N_{\mathrm{A}}\bm{I}_{N_{\mathrm{T}}}$ is satisfied, the calculation of the Moore-Penrose pseudo-inverse does not require explicit matrix inversion; instead, it reduces to a scaled transpose operation.:
This low-complexity variant of the QD-SMDS algorithm is hereafter referred to as QD-MRC-SMDS, and its pseudo-code is provided in Algorithm 3. Compared with the original QD-SMDS, QD-MRC-SMDS significantly reduces the overall complexity by eliminating expensive operations such as QSVD, matrix inversion, and Procrustes transformation. Instead, it enables direct coordinate estimation via simple quaternion-domain multiplications by utilizing a specific subset of the quaternion-domain GEK matrix.
The complexity reduction is also evident in the asymptotic sense. The most computationally intensive step is the matrix–vector multiplication $\bm{K}_{2}^{\mathsf{H}}\bm{\nu}_{\mathrm{AA}}$. Its asymptotic complexity scales as $O\left(N_{A}^{2}N_{\mathrm{T}}(N_{A}-1)/2\right)=O\left(N_{A}^{3}N_{\mathrm{T}}\right)$, i.e., linear in the number of targets $N_{\mathrm{T}}$.
@@ V-B Iterative QD-MRC-SMDS
Algorithm 4 Iterative QD-MRC-SMDS
1:Input:
3:The coordinates of all ANs.
4:Steps:
5:Construct $\bm{\nu}_{\mathrm{AA}}$, $\bm{K}_{2}$ and $\bm{K}_{3}$ via equations in Eqs. (28) and (32).
6:Initialize $\bm{\nu}_{\mathrm{AT}}$ using Eq. (51).
7:for $\tau=1,\ldots,\tau_{\mathrm{max}}$
8:Update $\bm{\nu}_{\mathrm{AT}}$ using Eq. (50).
9:end for
10:Obtain $\hat{\bm{\chi}}_{T}$ using Eq.(52).
@@
The computational complexity of iterative QD-MRC-SMDS is dominated by the multiplication of
$\begin{bmatrix}\bm{K}_{2}^{\mathsf{H}}&\bm{K}_{3}^{\mathsf{H}}\end{bmatrix}\in\mathbb{H}^{N_{A}N_{\mathrm{T}}\times M},$ (53)
with
$\begin{bmatrix}\bm{\nu}_{\mathrm{AA}}\\ \bm{\nu}_{\mathrm{AT}}^{(\tau)}\end{bmatrix}\in\mathbb{H}^{M\times 1},$ (54)
yielding a complexity of $O\bigl(N_{A}N_{\mathrm{T}}\bigl(N_{A}(N_{A}-1)/2+N_{A}N_{\mathrm{T}}\bigr)\bigr)=O\left(N_{A}^{3}N_{\mathrm{T}}+N_{A}^{2}N_{\mathrm{T}}^{2}\right)$, which adds an $O\left(N_{A}^{2}N_{\mathrm{T}}^{2}\right)$ term compared with standard QD-MRC-SMDS.
This iterative variant of the QD-MRC-SMDS algorithm is hereafter referred to as iterative QD-MRC-SMDS, and its pseudo-code is provided in Algorithm 4.